A 58kg boy runs at 6.5m/s grabs a hanging 10m long rope and swings out over the lake. He releases the rope when his velocity is zero.
a.What is the maximum tension in the rope during the swing?
1
Expert's answer
2020-05-08T16:21:30-0400
"T\cos{\theta}=mg"

"0.5mv^2=mgl(1-\cos{\theta})\to v^2=2gl\left(1-\frac{mg}{T}\right)"

"6.5^2=2(9.8)(10)\left(1-\frac{(58)(9.8)}{T}\right)"

"T=725\ N"